OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7 A. {(-1, 3), (0, 7/3), (2, 1)} B. {(3, -1), (7/3, 0), (1, 2)} C. {(5, -1), (7/2, 0), (2, 1)} D. {(-1, -3), (0, 7/3), (2, 1)}

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

since domain of x = (-1,0,2), the answer can be one out of A or D.(i.e. the x oval components) we also find that there are common options (0,7/3) and (2,1) Hence the decisive option is (-1,3) or (-1,-3) Now using (-1,3) in 2x+3y = 7 If 2.-1 + 3.3 = 7 i.e. -2 + 9 = 7 i.e. 7 = 7, which is true. Hence (-1,3) satisfies the equation. Answer - option A